# Simulations of Global Solar Convection with a Fully Compressible CHORUS++ Code

### v2

@@ Abstract
We present initial simulations of the solar convection zone using a fully compressible hydrodynamic CHORUS++ code and discuss preliminary analysis. Fluid dynamics simulation of the global solar convection is a critically important tool to access the dynamics of solar cycle variations. The CHORUS++ code robustly and efficiently solves the fully compressible hydrodynamic equations using a compact local spectral method and semi-unstructured grid system. Using CHORUS++, we simulate, for the first time, the solar convection shell from 0.7 to 0.99 of the solar radius, using the actual values of the total luminosity and the sidereal rotation rate. The simulation results include the longitudinally averaged rotation rate, reasonably agreeing with the observed solar-type differential rotation. The divergence of simulated mass flux infers that the anelastic-type models are appropriate for modeling the global solar convection, except for the outermost part of the Sun, for which the temporal scale of density variation is estimated at an order of days. The spherical harmonics analysis yields that the horizontal flows are dominant in the large-scale structure, and the degree of the anisotropy of the plasma flow is rather small and constant for the small-scale structures and for a wide range of the radius.
The Sun: corona, magnetic fields, solar wind
1 Introduction
Numerical simulations have offered several key assessments to the long-standing question of how plasma flows in the solar convection system influence and contribute to the 11-year solar-cycle variations of the magnetic field.
@@ 1 Introduction
For example, the anelastic hydrodynamics models (e.g., Miesch et al., 2000; Elliott et al., 2000) and those coupling with the magnetic field (e.g., Brun et al., 2004) successfully retrieved several stellar convection structures, such as the solar-like differential rotation (i.e. the faster rotations at low-latitude regions and slower rotations at the high-latitude regions) under some constraints, the meridional circulation (i.e. poleward flows near the surface and counterflows in the deep convection zone) and the laminar flow patterns when the stellar rotation rate is set high. The description of the mean-field MHD approach (e.g., Pipin & Kosovichev, 2024) offers a large-scale macroscopic view of azimuthally averaged flows and essential three-dimensional dynamo processes. The influence of the tachocline and the near-surface magnetic field has been studied by means of numerical implicit Large-Eddy Simulations (ILES) (e.g., Guerrero et al., 2016). The observation data analysis quantifies the near-surface plasma motions that manifest various forms of global flows, such as meridional flows and convective flows of different temporal and spatial scales (Rightmire-Upton et al., 2012; Hathaway et al., 2013). A notable novel approach is simulations of hydrodynamics (HD) and magnetohydrodynamics (MHD) with the sound-speed reduction method (e.g., Hotta et al., 2014, 2022), where the mass continuity equation is modified to reduce the sound speed and hence increase the simulation time step. This approach has proved efficient, in particular, in parallel computing because the simulation scheme can be highly localized and compact, while the basic nature of the MHD (or HD) is nearly fully taken into account.
@@
As shown in simulations of the interactions between the magnetic field flux and the plasma in the solar convection zone (e.g. Fan et al., 2003), compressibility may play a key role in the solar convection system. Despite limiting the scope to the HD case, density variations play primary roles in various processes, such as driving and maintaining the turbulent flows in the convection system, particularly in the near-surface rotational shear layer. The material cooled on the solar surface must be heavier because of the local pressure balance, and ultimately, the deviation becomes large enough to break the fragile local hydrostatic balance and start to flow downward. These nonlinear processes are treated correctly when and only when simulated with fully compressible models (e.g., Kitiashvili et al., 2023).
@@
Several factors have prevented the simulations from using the actual solar parameters. When the actual values of the solar luminosity and rotation rate are used, the evolution from the (usually semi-steady) initial state to the global convective flows will be slow. Hence, it is computationally expensive to simulate the solar convection zone with the actual solar parameters. A simulation that includes a near-surface region is challenging because of the very steep gradients of plasma density, temperature, and pressure. So far, it has only been done in local rectangular domains. It requires setting up small grid sizes in the radial direction to treat steep gradients, which requires us to adopt substantially smaller time steps in the ordinary explicit HD or MHD simulation model, leading to more computational costs.
@@
In part to mitigate these computational difficulties, many solar convection models have often used the so-called laminar benchmark case (Jones et al., 2011). A commonly used set of benchmark parameters includes large values of the stellar luminosity and the stellar rotation rate, which enhances the growth of convection and hence requires fewer computational resources. The assumption of incompressible flow (i.e., $\nabla\cdot(\varrho{\bf{V}})=0$) is often used for the same reasons. Although the simulations with the benchmark setting and/or the incompressible flows (or the reduced-sound-speed methods) successfully find several important aspects of the solar convection zone, it is desired to simulate the solar convection zone with a fully compressible code and with the actual solar parameters. In addition, if we conduct simulations with the compressible simulation, we can assess the appropriateness of the assumption of the divergence-free mass flux currently used in most solar-stellar global HD simulations.
@@
CHORUS++ (Chen et al., 2023) is an improved version of the CHORUS model (Compressible High-Order Unstructured Spectral differencing Wang et al., 2015, 2016). CHORUS has two notable features: the unstructured grid system and the spectral differencing method. CHORUS++ employs a cubed-sphere meshing algorithm (Ronchi et al., 1996) and improves CHORUS from iso-parametric mappings to transfinite mappings. As a result, CHORUS++ demonstrated sufficient numerical stability using the 6th- and 7th-order spectral difference methods Chen et al. (2023). The simulated volume is covered by six subdomains without any overlaps or gaps (e.g., Feng et al., 2010); hence, only flux points on a two-dimensional surface are needed to consider for exchanging physical information between the subdomains. The spectral differencing method is a subset of the flux reconstruction/correction procedure (e.g. Kopriva, 1994; Sun et al., 2007; Huynh, 2007, 2009). Each subdomain is discretized with semi-unstructured spherical grid assignment to avoid severe CFL (Courant–Friedrichs–Lewy) condition. The CHORUS++ code has been demonstrated to have improved computational efficiency over the original CHORUS code on distributed memory CPU systems in (Chen et al., 2023) to solve the benchmark cases of the stellar convection system robustly using a higher order spatial discretization. In this article, we use the CHORUS++ code with a few modifications in the simulation setting: The total stellar luminosity is set to the solar luminosity, the rotation rate is set to the solar mean sidereal rotation, and the inner and outer boundary surfaces are set at 0.70 $R_{\sun}$ and 0.99 $R_{\sun}$, respectively.
@@
This article is organized as follows. In Section 2, the CHORUS++ code and the simulation setting for this study are given. In Section 3, the simulation results are shown, and several properties of the derived plasma flows in the solar convection zone, such as the flow patterns and the divergence of the mass flux, are examined. In Section 4, the summary and conclusions are given.
2 Simulation Model
@@
CHORUS++ (Chen et al., 2023) uses the explicit five-stage and third-order Runge-Kutta method of stability preservation (SSPRK (5,3)) (Ruuth, 2005) to solve the set of equations of a fully compressible fluid in the rotating reference frame at a constant rotation rate (${\bf{\Omega_{o}}}$),
${\partial\varrho\over\partial t}$ $=$ $-\nabla(\varrho{\bf{U}}),$ (1)
@@
${\partial(\varrho{\bf{U}})\over\partial t}$ $=$ $-\nabla\left(\varrho{\bf{U}}:{\bf{U}}+P{\cal I}\right)+\nabla\left(\bar{\bf\tau}\right)+\varrho{\bf{g}}-2\varrho{\bf{\Omega_{o}}}\times{\bf{U}},$ (2)
${\partial{\cal E}\over\partial t}$ $=$ $-\nabla\cdot\left({\bf{U}}({\cal E}+P)\right)+\nabla\left({\bf{U}}\cdot\bar{\bf\tau}-\bar{{\bf q}}\right)+\varrho{\bf{g}}\cdot{\bf{U}},$ (3)
where $\varrho$, ${\bf{U}}$, and ${\cal E}$ are the density, velocity vector, and total energy. Under the ideal gas assumption, the gas pressure $P$ is given as $P=\varrho{\cal R}T$ with the gas constant ${\cal R}$, and the total energy is given as ${\cal E}=\frac{P}{\gamma-1}+{1\over 2}\varrho U^{2}$, with the specific heat ratio $\gamma(=5/3)$. As in Chen et al. (2023), the centrifugal force is not included in this study.
@@
$\bar{\bf\tau}=\mu\left(\nabla{\bf{U}}+(\nabla{\bf{U}})^{T}+\lambda(\nabla\cdot{\bf{U}}){\cal I}\right)$ (4)
with the dynamic viscosity, $\mu=\nu\varrho$ ($\nu$ is the kinematic viscosity coefficient), and $\lambda=-(2/3)\mu$. The heat flux ${\bf b}$ is given as
${\bf b}=-\kappa\varrho T\nabla S-\kappa_{r}\varrho C_{p}\nabla T$ (5)
@@
with the specific heat at constant pressure $C_{p}$. The extra terms are due to the gravity of the Sun and the Coriolis force.
In the CHORUS++ code, the velocity vector ${\bf{U}}$ is defined in the Cartesian system, ${\bf{U}}=(u,v,w)$ or ${\bf{U}}=(V_{x},V_{y},V_{z})$, and the numerical fluxes normal to the cell faces are calculated through the Bassi-Rebay I scheme Bassi & Rebay (1997). The governing equations in the physical domain are transformed into local coordinates of a computational cell using the transfinite mapping approach. The transfinite mapping approach is a major improvement over the original CHORUS code Wang et al. (2015) because it can represent the curves of the edges and faces of each physical element within the machine precision (Chen et al., 2023). For the initial study, we set the grid size in the radial direction at about 0.003 $R_{\sun}$ and the heliocentric angles of the numerical cells at 0.8 $\sim$ 0.95 degrees.
2.1 Model parameters, initial values, and simulation setup
We used the same parameters as in the benchmark test case (Chen et al., 2023) except the stellar luminosity, rotation rate, and radius of the upper boundary sphere surface. In conjunction with the change of the radius of the top boundary surface ($r=R_{\rm top}$), we need to adjust the plasma density there. The parameters set in the presented study are tabulated in Table 1. For reference, the benchmark parameters and some of the actual solar properties are also given in this table.
Table 1: Parameters, constants, and representative numbers for (a) the actual Sun, (b) the presented cases, and (c) the benchmark case. The numbers for the benchmark case are from Chen et al. (2023) (and/or Wang et al. (2015)). The constants are set as follows: Gravitational constant $G=6.67\cdot 10^{-8}$ (${\rm cm}^{3}{\rm g}^{-1}{\rm s}^{-2}$), the solar mass $M_{\sun}=1.98891\cdot 10^{33}$ (g), the specific heat ratio $\gamma=5/3$, the gas constant ${\cal R}=1.4\cdot 10^{8}$ (erg ${\rm g}^{-1}{\rm K}^{-1}$, the specific heat at constant pressure $C_{P}=3.5\cdot 10^{8}$ (erg ${\rm g}^{-1}{\rm K}^{-1}$), the solar luminosity $L_{\sun}=3.846\cdot 10^{33}$ (erg s-1), and the solar sidereal rotation rate $\Omega_{\sun}=3.0\cdot 10^{-6}$ (rad s-1). The parameter $D$ is the thickness of the convection zone ($D=R_{\rm top}-R_{\rm bot}$).
(a) the Sun (b) presented case (c) benchmark

### v3

@@ Abstract
Fluid-dynamics simulations of global solar convection are a critically important tool for assessing the dynamics of the solar interior. However, simulation studies with a fully compressible hydrodynamics code are not yet common. The CHORUS++ code solves robustly and efficiently the fully compressible hydrodynamics equations using a compact local spectral method and semi-unstructured grid system. Using the CHORUS++ code, we simulate the solar interior plasma flows from 0.7 to 0.99 of the solar radius using the values of the solar total luminosity and the sidereal rotation rate. In this paper, we analyze the simulated global flow structures before the statistically stable state to assess the compressibility of the plasma flows obtained with the fully compressible hydrodynamic code. The divergence of mass flux with the compressible model is overall small in the examined state before reaching the equilibrium state of the convection, which implies that the differences between the fully compressible flows and those obtained with the anelastic, incompressible, or linear-equation models are small in the simulated inner part of the convection zone. Although a fully-relaxed stationary convection has not been achieved yet in the examined state, the model qualitatively reproduces the solar-type differential rotation. Simulations for longer periods are needed to achieve the system relaxation state. This work, assessing the early phase of the hydrodynamic evolution of solar convection, is our first step toward a better understanding of the nature of solar convection and the dynamo processes.
I Introduction
Helioseismic observations from space using the Michelson Doppler Imager (MDI) on board the Solar and Heliospheric Observatory (SOHO) (Scherrer et al., 1995) and the Helioseismic and Magnetic Imager (HMI) on board the Solar Dynamics Observatory (SDO) (Scherrer et al., 2012), as well as from the ground-based Global Oscillations Network Group (GONG) (Harvey et al., 1996), revealed the solar internal rotation and its variations with the solar cycle in great detail. It has been established that the latitudinal differential rotation known from observations of the solar surface extends through the whole convection zone, with two distinct regions, characterized by high radial gradients of the rotation rate, the tachocline at the bottom, and the near-surface shear layer at the top. In the bulk of the convection zone, the solar rotation rate has a complicated radial and latitudinal structure with iso-rotational lines forming conical surfaces (Schou et al., 1998). The observations also revealed cyclic variations in the solar rotation, creating a system of zonal flows (‘torsional oscillations’), migrating from mid-latitudes towards the equator and the polar regions (Howe et al., 2000). In addition, local helioseismology methods provided measurements of the meridional circulation and properties of subsurface convection (for a review of recent advances in global and local helioseismology, see Kosovichev et al. (2025) and references therein).
@@ I Introduction
The observed properties of the solar internal rotation and its variations have not been fully reproduced by theoretical models. Nevertheless, numerical simulations have provided several key insights into the long-standing question of how plasma flows in the solar convection system interact with rotation and contribute to the 11-year solar-cycle variations of the magnetic field. For example, the anelastic hydrodynamics models (e.g., Miesch et al., 2000; Elliott et al., 2000) and those coupling with the magnetic field (e.g., Brun et al., 2004) successfully retrieved several stellar convection structures, such as the solar-like differential rotation (i.e. the faster rotations at low-latitude regions and slower rotations at the high-latitude regions) under some constraints, the meridional circulation (i.e. poleward flows near the surface and counterflows in the deep convection zone) and the laminar flow patterns when the stellar rotation rate is set high. The description of the mean-field MHD approach (e.g., Pipin and Kosovichev, 2024) offers a large-scale macroscopic view of azimuthally averaged flows and essential three-dimensional dynamo processes. The influence of the tachocline and the near-surface magnetic field has been studied by means of numerical implicit Large-Eddy Simulations (ILES) (e.g., Guerrero et al., 2016). The observation data analysis quantifies the near-surface plasma motions that manifest various forms of global flows, such as meridional flows and convective flows of different temporal and spatial scales (Rightmire-Upton et al., 2012; Hathaway et al., 2013). A notable novel approach is simulations of hydrodynamics (HD) and magnetohydrodynamics (MHD) with the sound-speed reduction method (e.g., Hotta et al., 2014, 2022), where the mass continuity equation is modified to reduce the sound speed and hence increase the simulation time step. This approach has proved efficient, in particular, in parallel computing because the simulation scheme can be highly localized and compact, while the basic nature of the MHD (or HD), such as the time-dependent nonlinear behavior of the hyperbolic equation system in multi-dimensional space, and the conservation of mass, momentum, and energy, is nearly fully taken into account.
@@
As shown in simulations of the interactions between the magnetic field flux and the plasma in the solar convection zone (e.g. Fan et al., 2003), compressibility may play a key role in the solar convection system. Despite limiting the scope to the HD case, density variations play a primary role in various processes, such as driving and maintaining the turbulent flows in the convection system, particularly in the near-surface rotational shear layer. The material cooled on the solar surface must be heavier because of the local pressure balance, and ultimately, the deviation becomes large enough to break the fragile local hydrostatic balance and start to flow downward. These nonlinear processes are treated correctly when and only when simulated with fully compressible models like those for near-surface models (e.g., Kitiashvili et al., 2023).
@@
Several factors have prevented the simulations from using the actual solar parameters. When the actual values of the solar luminosity and rotation rate are used, the evolution from the (usually semi-steady) initial state to the global convective flows will be slow. Hence, it is computationally expensive to simulate the solar convection zone with the actual solar parameters. A simulation that includes a near-surface region is challenging because of the very steep gradients of plasma density, temperature, and pressure. So far, it has only been done in local rectangular domains. It requires setting up small grid sizes in the radial direction to treat steep gradients, which requires us to adopt substantially smaller time steps in the ordinary explicit HD or MHD simulation model, leading to increased computational costs.
@@
In part to mitigate these computational difficulties, a simulation model for solar convection had to apply a different set of assumptions and simulation parameters to define the simulated situations. A widely used set of parameters includes larger values of the stellar luminosity and/or the rotation rate than the actual solar parameters (e.g., Wang et al., 2015), which enhances the growth of convection and hence requires fewer computational resources. Besides the parameter modification, the assumption of incompressible flow (i.e., $\nabla\cdot(\varrho\vec{V})=0$) or the anelastic model (Gough, 1969) are often used, in part, to achieve computational stability. The simulations with these assumptions successfully find several important aspects of the solar convection zone. Nonetheless, to assess the difference and/or similarity between the fully compressible flows and the anelastic or incompressible flows, it is critically important to conduct the simulation of the global solar convection zone with a fully compressible code and with the actual solar parameters.
@@
CHORUS++ (Chen et al., 2023) is an improved version of the CHORUS model (Compressible High-Order Unstructured Spectral differencing Wang et al., 2015, 2016). CHORUS has two notable features: the unstructured grid system and the spectral differencing method. CHORUS++ employs a cubed-sphere meshing algorithm (Ronchi et al., 1996) and improves CHORUS by replacing iso-parametric mappings with transfinite mappings. As a result, CHORUS++ demonstrated sufficient numerical stability using the 6th- and 7th-order spectral difference methods (Chen et al., 2023). The simulated volume is covered by six subdomains without any overlaps or gaps (e.g., Feng et al., 2010); hence, only flux points on a two-dimensional surface are needed to consider for exchanging physical information between the subdomains. The spectral differencing method is a subset of the flux reconstruction/correction procedure (e.g. Kopriva, 1994; Sun et al., 2007; Huynh, 2007, 2009). Each subdomain is discretized with a semi-unstructured spherical grid assignment to avoid severe CFL (Courant-Friedrichs-Lewy) condition. The CHORUS++ code has been demonstrated to achieve improved computational efficiency over the original CHORUS code on distributed-memory CPU systems (Chen et al., 2023) to solve the stellar convection system robustly using a higher-order spatial discretization.
In this article, we use the CHORUS++ code with a few modifications in the simulation setting: the total stellar luminosity is set to the solar luminosity, the rotation rate is set to the solar mean sidereal rotation, and the inner and outer boundary surfaces are set at 0.70 $R_{\sun}$ and 0.99 $R_{\sun}$, respectively. In particular, by setting the outer boundary surface at 0.99 $R_{\sun}$, we can include the regions with very steep radial gradients of the plasma temperature and density. With the outer boundary surface close to 1.0 $R_{\sun}$, the simulations can consider the regions with wider ranges of plasma quantities (i.e., the contrast of the plasma values at the bottom and top boundaries), and we can assess the differences and similarities between a fully compressible flow system and non-fully-compressible systems.
Ideally, the analysis of the simulation results would be conducted for the states of the statistically stable convective system after the onset of the convection and an overshoot saturation of the total kinetic energy. However, the computation, even with the CHORUS++, requires a substantial amount of computation when the top boundary surface is set close to 1 $R_{\sun}$, because the grid points in the radial direction must be assigned with a smaller grid size to capture and handle the steep gradients of the plasma quantities at $r\lesssim 1R_{\sun}$. Therefore, in this article, we use the simulated data before reaching the statistically stable convection state. Instead of examining the statistically stable convective system, the analysis is focused on the convection structures in the early phase of convection, in which the net convective kinetic energy grows exponentially in time, saturates, and starts seeking the stable state. The analyzed quantities here include the divergence of the mass fluxes to examine the degree of compressibility in the fully compressible system.
@@
This article is organized as follows. In Section II, the CHORUS++ code and the simulation setting for this study are given. In Section III, the simulation results are shown, and several properties of the derived plasma flows in the solar convection zone, such as the flow patterns and the divergence of the mass flux, are examined. In Section IV, the summary and conclusions are given.
II Simulation Model
@@
CHORUS++ (Chen et al., 2023) uses the explicit five-stage and third-order Runge-Kutta method of stability preservation (SSPRK (5,3)) (Ruuth, 2005) to solve the set of equations of a fully compressible fluid in the rotating reference frame at a constant rotation rate ($\vec{\Omega_{o}}$),
${\partial\varrho\over\partial t}$ $=$ $-\nabla(\varrho\vec{U}),$ (1)
@@
${\partial(\varrho\vec{U})\over\partial t}$ $=$ $-\nabla\left(\varrho\vec{U}\otimes\vec{U}+P{\cal I}\right)+\nabla\left(\bar{\bf\tau}\right)+\varrho\vec{g}-2\varrho\vec{\Omega_{o}}\times\vec{U},$ (2)
${\partial{\cal E}\over\partial t}$ $=$ $-\nabla\cdot\left(\vec{U}({\cal E}+P)\right)+\nabla\left(\vec{U}\cdot\bar{\bf\tau}-\bar{{\bf q}}\right)+\varrho\vec{g}\cdot\vec{U},$ (3)
[…]
@@
$\bar{\bf\tau}=\mu\left(\nabla\vec{U}+(\nabla\vec{U})^{T}+\lambda(\nabla\cdot\vec{U}){\cal I}\right)$ (4)
with the dynamic viscosity $\mu=\nu\varrho$, where $\nu$ is the kinematic viscosity coefficient, and $\lambda=-(2/3)\mu$. The heat flux $\bf q$} is given as
${\bf q}=-\kappa\varrho T\nabla S-\kappa_{r}\varrho C_{p}\nabla T$ (5)
@@
with the specific heat at constant pressure $C_{p}$. The extra terms, $\varrho\vec{g}$ and $2\varrho\vec{\Omega_{o}}\times\vec{U}$, are for the gravity and the Coriolis force, respectively.
The impenetrable plasma flow, $V_{r}=\vec{U}\cdot\hat{r}=0$ ($\hat{r}$ is the unit vector of the heliocentric position vector $\vec{r}$, $\hat{r}:=\vec{r}/|\vec{r}|$, or the outward-pointing normal unit vector to the boundary surface), is applied to the top and bottom boundary surfaces. In addition, the stress-free condition is achieved by resetting the shear stress components associated with the normal direction to the boundary sphere ($\tau_{r,\theta}$, $\tau_{r,\phi}$, $\tau_{\theta,r}$, and $\tau_{\phi,r}=0$ when writing the terms in the spherical coordinate system) to zero. On the bottom boundary surface, the heat flux $\bf q$} is fixed at $L_{\sun}/(4\pi R_{\rm bot}^{2})\hat{r}$.
In the CHORUS++ code, the velocity vector $\vec{U}$ is defined in the Cartesian system, $\vec{U}=(u,v,w)$ or $\vec{U}=(V_{x},V_{y},V_{z})$, and the numerical fluxes normal to the cell faces are calculated through the Bassi-Rebay I scheme (Bassi and Rebay, 1997). The governing equations in the physical domain are transformed into local coordinates of a computational cell using the transfinite mapping approach. The transfinite mapping approach is a major improvement over the original CHORUS code (Wang et al., 2015) because it can represent the curves of the edges and faces of each physical element within the machine precision (Chen et al., 2023). In the presented study, we set the grid size in the radial direction at about 0.001563 $R_{\sun}$ and the heliocentric angles of the numerical cells at 0.4 $\sim$ 0.5 degrees.
II.1 Model parameters, initial values, and simulation setup
We used the same method as in Chen et al. (2023) to set up the initial values, except that the stellar luminosity, rotation rate, the radius of the upper boundary sphere surface, and the initial plasma density at the upper boundary surface are altered. As in Chen et al. (2023), the initial radial profiles of the plasma quantities are generated in accordance with the so-called laminar benchmark (Jones et al., 2011). In conjunction with the change of the radius of the top boundary surface ($r=R_{\rm top}$), we need to adjust the plasma density there. The parameters set in the presented study are tabulated in Table 1. For reference, the parameters used in the earlier work using the CHORUS++ code (Chen et al., 2023) and some of the actual solar properties are also given in this table.
Table 1: Parameters, constants, and representative numbers for (a) the actual Sun, (b) the presented case, and (c) the numbers used in Chen et al. (2023) (and/or Wang et al. (2015)). The constants are set as follows: Gravitational constant $G=6.67\cdot 10^{-8}$ (${\rm cm}^{3}{\rm g}^{-1}{\rm s}^{-2}$), the solar mass $M_{\sun}=1.98891\cdot 10^{33}$ (g), the specific heat ratio $\gamma=5/3$, the gas constant ${\cal R}=1.4\cdot 10^{8}$ (erg ${\rm g}^{-1}{\rm K}^{-1}$, the specific heat at constant pressure $C_{P}=3.5\cdot 10^{8}$ (erg ${\rm g}^{-1}{\rm K}^{-1}$), the solar luminosity $L_{\sun}=3.846\cdot 10^{33}$ (erg s-1), and the solar sidereal rotation rate $\Omega_{\sun}=3.0\cdot 10^{-6}$ (rad s-1). The parameter $D$ is the thickness of the convection zone ($D=R_{\rm top}-R_{\rm bot}$).
(a) the Sun (b) presented case (c) Chen et al. (2023)